A farmer sold bushels of wheat and bushels of oats. For 425 bushels, he received $481.25. How many bushels of each did he sell if he received $2.08 per bushel of wheat and $.94 per bushels of oat?
Answer by Boreal(15235)

W+O=425
2.08W+0.94O=481.25
multiply top by -0.94 to eliminate O
-0.94W-0.94O=399.5
1.14W=81.75
W=71.71 bushels wheat
O =353.29 bushels oats
This seems odd as I would expect integers.
71.71*2.08=$149.16
353.29*0.94=$332.09
They check.